# State-space models are accurate and efficient neural operators for dynamical systems

### v1

@@ 4.5 Extrapolation to Long-Time Integration Problems
Here are the implementation details. We generate 10K train and 10K test data with a sensor/discretization size of $100T$ where $T$ is the terminal time, i.e., the sensors are on $\{0.01,0.02,\cdots,T-0.01,T\}$ where the input function $u(t)$ and the target function $s(t)/s_{1}(t)$ are discretized. The training loss is MSE, while the test metric is the relative $L_{2}$ error to ensure comparability between tests on various intervals. We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [26] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training. We keep the number of parameters of different models similar for a fair comparison, following the first experiment exactly.
@@
Test Interval GRU LSTM Transformer Oformer-V Oformer-G Oformer-F GNOT Mamba
[0,1] 6.813E-4 7.707E-4 1.069E-2 5.071E-4 3.905E-4 3.856E-4 7.195E-4 2.310E-4
[0,2] 9.833E-2 1.715E-1 8.262E-1 8.687E-1 8.589E-1 8.658E-1 7.737E-1 3.060E-2
[0,3] 2.827E-1 4.697E-1 8.962E-1 9.182E-1 9.279E-1 9.682E-1 9.486E-1 1.550E-1
[0,4] 4.589E-1 7.022E-1 9.385E-1 9.529E-1 9.600E-1 9.971E-1 9.693E-1 3.570E-1
@@ 4.6 Chaotic Lorenz System with Various Initial Conditions
So far, we have focused on solving sequence-to-sequence problems in dynamical systems. This is challenging since the model’s input sequence is a high-dimensional vector due to the long time plus some additional ODE coefficient/parameter. Another crucial setting is to solve ODE and dynamical systems with initial conditions only, i.e., the input is just the initial conditions, and the output is the ODE solution over a long time. For the test case we consider a chaotic Lorenz system with random initial conditions following Li et al. [29]:
@@
We aim to learn the operator from the random initial condition to the state variable $x(t)$. We opt for $\sigma=10$ and $\beta=8/3$ and $\rho=28$. More specifically, we generate 10K data with the initial condition drawn from $x(0)\sim\text{Unif}(-10,10),y(0)\sim\text{Unif}(-20,30),z(0)\sim\text{Unif}(10,40)$ and 90% for training while another 10% for testing. The target is the ODE solution over $[0,T]$ discretized with a step size 1e-3, i.e., we wish to predict the ODE solution values on the $1000\cdot T$ grid points $\{0.001,0.002,\cdots,T\}$. We test for $T=1,2,3$. The training loss is MSE, while the test metric is the relative $L_{2}$ error to ensure comparability between tests on various terminal times. We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [26] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training.
Regarding tested models, since the problem is a long sequence, we test the efficient LSTM, GRU, and Mamba with two different structures and model sizes. For LSTM and GRU, we test one-layer models with 32 or 64 hidden units. For Mamba, we test a one-block model with 16 or 32 model dimensions and intermediate dimensions.
GRU1 GRU2 LSTM1 LSTM2 Mamba1 Mamba2
Layer Num 1 1 1 1 1 1
Hidden Size 32 64 32 64 16 32
Params 6529 25345 8641 33665 9679 22593
$T=1$ Memory 2397MB 3039MB 2507MB 3257MB 2157MB 2265MB
Time 59min 75min 58min 61min 95min 116min
Rel. $L_{2}$ Err 3.909E-2 NaN 3.720E-2 6.158E-1 3.991E-2 3.417E-2
$T=2$ Memory 3229MB 4511MB 3447MB 4949MB 2919MB 3143MB
Time 117min 146min 94min 114min 193min 230min
Rel. $L_{2}$ Err NaN NaN NaN NaN 1.338E-1 1.230E-1
$T=3$ Memory 4063MB 5983MB 4389MB 6637MB 3687MB 4033MB
Time 181min 215min 134min 169min 296min 346min
Rel. $L_{2}$ Err NaN NaN NaN NaN 2.470E-1 2.157E-1
@@
Results are shown in Table 8. In the chaotic Lorenz system with a terminal time equal to $T=2$ or $T=3$, we find that LSTM and GRU models of various sizes diverge to not-a-number (NaN) values during training due to their model instability. Chaotic systems are known for their instability and sensitivity to minor changes in their initial values. A minor difference in initial values will lead to a considerable difference in ODE trajectories, especially during long-time integration. This, in turn, implies that the model has to be able to capture the subtle differences in initial values and react differently, which is a challenging task. Hence, we observe that LSTM and GRU converge well in the $T=1$ case where the time is short and the chaotic dynamics has not developed yet. However, as we reach longer times, LSTM and GRU diverge due to the chaotic dynamics even though we train them using a decaying learning rate for stabilization. This reflects RNNs’ inherent instability in lengthy sequence modeling (length = 1K, 2K, and 3K here) due to unstable gradient explosion even for its modern variants, namely LSTM and GRU [2, 38]. In contrast, Mamba always converges in all systems, demonstrating its superior stability and robustness regardless of the long time integration.
@@ 4.7 Quantifying the Extrapolation Error
We have extensively tested Mamba’s extraordinary out-of-distribution (OOD) extrapolation capability in (1) OOD operator learning following Cao et al. [4] in Section 4.3, (2) extrapolation to long-time interation problems while only training on shorter time 4.5, (3) extrapolation to treatment schedule parameters in the PK-PD model in Section 4.8.4. In this section, we follow Zhu et al. [55] to quantify further various models’ extrapolation errors under different train and test input function distributions.
Regarding the ODE operators, we test the antiderivative and the gravity pendulum operators in Section 4.1 and keep all settings, e.g., implementation details, training procedures, hyperparameters, and model structures, the same as those in Section 4.1 except the train and test input function distributions.
Following Zhu et al. [55], the input functions $u(t)$ are randomly drawn from a mean zero Gaussian random field (GRF) with a length scale parameter $l$, i.e., the input functions follow the distribution $\mathcal{GP}\left(0,k_{l}(x,y)\right)$ where $k_{l}(x,y)=\exp\left(-\frac{(x-y)^{2}}{2l^{2}}\right)$ is the RBF kernel with length scale $l$ and $\mathcal{GP}$ denotes a Gaussian process. A smaller correlation length scale $l$ leads to less smooth functions, while a larger correlation length scale $l$ leads to smoother functions.
@@
The detailed quantitative results are shown in Table 12 (gravity pendulum Ex-), Table 13 (gravity pendulum Ex+), Table 14 (anti-derivative Ex-), Table 15 (anti-derivative Ex+) in Appendix B. Their visualization is shown in Figure 9 (gravity pendulum Ex-), Figure 10 (gravity pendulum Ex+), Figure 11 (anti-derivative Ex-), Figure 12 (anti-derivative Ex+). Each figure shows the visualized results for one of the four settings, gravity pendulum or anti-derivative operator in the Ex- or the Ex+ setting. Furthermore, in each figure, we show Mamba’s comparison with RNNs, including GRU and LSTM, in subfigure (a), with neural operators (DeepONet, FNO, and LNO) in (b), and with transformers (Transformer, Oformers, and GNOT) in (c).
Regarding the pendulum problem, Mamba is the best in both Ex- problems (Figure 9 and Table 12) and Ex+ (Figure 10 and Table 13). More concretely, in the Ex- setting (Figure 9 and Table 12), Mamba’s error remains stable under the test distribution shift. Although certain transformers (Oformer-G and Oformer-F) also exhibit stability against test distribution shift, their error is larger than Mamba. Other models, including RNNs and neural operators, exhibit increases in errors when the test distribution shift is more significant, i.e., their error grows larger as $l_{test}=0.1$ gradually moves to $l_{test}=1$. Meanwhile, regarding the Ex+ setting of gravity pendulum (Figure 10 and Table 13), all models’ errors grow with test distribution shift as $l_{test}=1$ gradually moves to $l_{test}=0.1$ where the training set length scale is $l_{train}=1$. However, Mamba exhibits the slowest error growth and consistently performs the best.
Regarding the anti-derivative operator, Mamba is the best in the Ex- setting (Figure 11 and Table 14), and its Ex- error remains stable. We also observe certain transformers (Oformer-G and Oformer-F) can exhibit stability against test distribution shift but perform worse than Mamba in general. Moreover, other models, including RNNs and neural operators, exhibit increases in errors when the test distribution shift is more significant, i.e., their error grows larger as $l_{test}=0.1$ gradually moves to $l_{test}=1$. This is similar to the Ex- case in the gravity pendulum case. Regarding the anti-derivative operator’s Ex+ setting (Figure 12 and Table 15), Mamba is good in the majority of test cases. Mamba is only beaten in the case with $l_{train}=1$ and $l_{test}=0.1$, although not by much and it is close to the best model. Quantitatively, Mamba’s relative $L_{2}$ error for the anti-derivative operator’s Ex+ setting with $l_{train}=1$ and $l_{test}=0.1$ is 3.161E-2, while the best model LSTM’s is 2.204E-2.
@@ 4.8.1 Model and Data Generation
Simeoni et al. [47] introduced a streamlined PK-PD model to simplify the complex dynamics of tumor growth, making it more applicable for the preclinical development of oncology drugs. This model, formulated with a system of ODEs, connects the drug dosing regimen to the tumor growth dynamics observed in animal studies. The model differentiates between untreated and treated animals; in untreated animals, tumor growth follows an initial exponential phase and then a linear phase, whereas in treated animals, the growth rate decreases in relation to the drug concentration and the number of proliferating tumor cells. The differential equation for the unperturbed growth of the tumor mass $w(t)$ is given by:
@@
In this work, we aim to learn the mapping from $I_{1}(t)$ in PK model, representing the drug dose and administration schedule in a multi-dose regimen, to the patient-specific function $w(t)$, which reflects tumor size based on individual patient PD model parameters $k_{1}$ and $k_{2}$. These parameters indicate disparate responses to treatment across different species.
@@ 4.8.2 Learning the PK-PD Model’s Operator
We generated a dataset containing 5000 species for both the PK and PD models based on the procedure in Section 4.8.1. The dataset comprises the dose amount and dose schedule as inputs (represented as $I_{1}(t)$ in the PK model), the species-specific PD model parameters $k_{1}$ and $k_{2}$, as well as $w(t)$, which is the model output. Upon feeding the model with the values of $k_{1}$, $k_{2}$, and $I_{1}(t)$, the resulting output is the tumor size in grams. It is observed that different dynamics and tumor sizes are exhibited by different species or treatment schedules.
@@
Here are the implementation details. The training loss is MSE, and the test metric is the relative $L_{2}$ error. We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [26] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training. We keep the number of parameters of different models similar for a fair comparison. We run the code for five independent random seeds and report the average.
@@
The model structure hyperparameter settings are summarized below. LSTM and GRU are one-layer models with 32 hidden dimensions. DeepONet’s trunk and branch nets are all five-layer networks with 24 hidden dimensions. FNO is a three-layer model with 16 hidden dimensions and 8 modes. We choose a one-layer LNO with 8 hidden dimensions and 4 modes and another one-layer LNO with 16 hidden dimensions and 8 modes. We also tried larger LNOs, but this small LNO model performs the best on the validation set. We report both LNO performances We opt for a four-layer Transformer with 12 hidden dimensions, 48 feedforward dimensions, and 4 heads. For all three Oformers with vanilla, Galerkin, or Fourier attention, we opt for three-layer Oformers with 12 hidden dimensions, 12 feedforward dimensions, and 4 heads. GNOT has three layers, 8 hidden dimensions and 8 feedforward dimensions, 4 heads, and 2 experts. Regarding Mamba, we opt for a block with skip connection and layer normalization, with 16 model dimensions and 16 intermediate dimensions.
The PK-PD Model
Model Params Memory Time Rel. $L_{2}$ Err Model Params Memory Time Rel. $L_{2}$ Err
LSTM 8641 1985MB 16min 2.117E-3 Transformer 7633 4287MB 170min 1.250E-2
GRU 6529 1925MB 16min 2.125E-3 Oformer-V 8569 3131MB 119min 6.246E-3
DeepONet 38568 1425MB 15min 4.732E-1 Oformer-G 8497 1603MB 110min 1.107E-2
FNO 7057 2019MB 26min 6.544E-3 Oformer-F 8497 1603MB 110min 8.649E-3
LNO (Large) 6769 1991MB 24min 8.269E-1 GNOT 8695 1711MB 234min 4.115E-3
LNO (Small) 1913 1935MB 22min 5.295E-3 Mamba 9697 1707MB 19min 1.685E-3
@@
Table 9: Results for the PK-PD model corresponding to Section 4.8.2. The best model is bold, the second best is red, and the third best is blue. We visualize some inputs and Mamba’s predictions in Figure 13. We visualize Mamba’s relative $L_{2}$ error concerning time in Figure 14.
@@
Figure 13: Visualization of Mamba’s prediction in PK-PD model corresponding to Section 4.8.2. We plot the input values and the outputs for test data. (a) and (b): input/output (io) for test data No. 0; (c) and (d): io for test data No. 100; (e) and (f): io for test data No. 200; (g) and (h): io for test data No. 300. Note that the heights and positions of the input spikes vary for different data sets, and the PD parameters differ based on each individual’s response to treatment. Moreover, we visualize Mamba’s relative $L_{2}$ error with respect to time in Figure 14. The full quantitative results are shown in Table 9.
@@
Results are shown in Table 9. Mamba continues to be the best-performing model among multiple strong baselines. We visualize the inputs and outputs of this PK-PD model in Figure 13 and plot Mamba’s relative $L_{2}$ error with respect to time in Figure 14. The results show that Mamba can capture the PK-PD model’s operator containing discontinuous inputs. The Mamba model’s relative $L_{2}$ error of 1.685E-3 is equivalent to an error of 0.1685% of the tumor size. Thus, for a tumor volume of 10 grams, the resulting error is 16.85 milligrams.

### v2

@@ 4.5 Extrapolation to Long-Time Integration Problems
Here are the implementation details. We generate 10K train and 10K test data with a sensor/discretization size of $100T$ where $T$ is the terminal time, i.e., the sensors are on $\{0.01,0.02,\cdots,T-0.01,T\}$ where the input function $u(t)$ and the target function $s(t)/s_{1}(t)$ are discretized. The training loss is MSE, while the test metric is the relative $L_{2}$ error to ensure comparability between tests on various intervals. We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [27] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training. We keep the number of parameters and model structures of different models similar for a fair comparison, following the first experiment (Section 4.1) exactly since the training dataset and the validation setup are the same.
@@
Test Interval GRU LSTM Transformer Oformer-V
[0,1] 8.292E-4$\pm$3.426E-4 7.718E-4$\pm$2.517E-4 2.207E-2$\pm$2.697E-4 4.324E-4$\pm$2.897E-5
[0,2] 9.169E-2$\pm$1.545E-2 1.596E-1$\pm$2.144E-2 9.190E-1$\pm$1.702E-2 8.118E-1$\pm$4.582E-2
[0,3] 2.669E-1$\pm$3.573E-2 4.302E-1$\pm$6.090E-2 9.625E-1$\pm$9.573E-3 9.452E-1$\pm$1.244E-2
[0,4] 4.411E-1$\pm$4.031E-2 6.161E-1$\pm$8.320E-2 9.747E-1$\pm$3.231E-3 9.638E-1$\pm$1.075E-2
Test Interval Oformer-G Oformer-F GNOT Mamba
[0,1] 4.228E-4$\pm$1.457E-4 4.096E-4$\pm$1.479E-4 3.700E-4$\pm$8.855E-6 2.175E-4$\pm$1.533E-5
[0,2] 8.145E-1$\pm$5.908E-2 8.165E-1$\pm$1.849E-3 7.099E-1$\pm$1.626E-2 2.823E-2$\pm$5.125E-3
[0,3] 9.473E-1$\pm$1.444E-2 9.460E-1$\pm$4.959E-3 9.120E-1$\pm$1.538E-2 1.475E-1$\pm$1.431E-2
[0,4] 9.622E-1$\pm$4.119E-3 9.522E-1$\pm$8.819E-3 9.515E-1$\pm$1.623E-2 3.451E-1$\pm$1.952E-2
@@ 4.6 Chaotic Lorenz System with Various Initial Conditions
So far, we have focused on solving sequence-to-sequence problems in dynamical systems. This is challenging since the model’s input sequence is a high-dimensional vector due to the long time plus some additional ODE coefficient/parameter. Another crucial setting is to solve ODE and dynamical systems with initial conditions only, i.e., the input is just the initial conditions, and the output is the ODE solution over a long time. For the test case, we consider a chaotic Lorenz system with random initial conditions following Li et al. [30]:
@@
We aim to learn the operator from the random initial condition to the state variable $x(t)$. We opt for $\sigma=10$ and $\beta=8/3$ and $\rho=28$. More specifically, we generate 10K data with the initial condition drawn from $x(0)\sim\text{Unif}(-10,10),y(0)\sim\text{Unif}(-20,30),z(0)\sim\text{Unif}(10,40)$ and 90% for training while another 10% for testing. The target is the ODE solution over $[0,T]$ discretized with a step size 1e-3, i.e., we wish to predict the ODE solution values on the $1000\cdot T$ grid points $\{0.001,0.002,\cdots,T\}$. We test for $T=1,2,3$. The training loss is MSE, while the test metric is the relative $L_{2}$ error to ensure comparability between tests on various terminal times. We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [27] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training. We append the temporal grid to the input data tensor for augmentation.
Regarding tested models, since the problem is a long sequence, we test the efficient LSTM, GRU, and Mamba with two different structures and model sizes. For LSTM and GRU, we test one-layer models with 32 or 64 hidden units. For Mamba, we test a one-block model with 16 or 32 model dimensions and intermediate dimensions. The goal is to test models with similar and comparable structures (all of them are one-layer models) but various hidden sizes. LSTM and GRU with 32 (64) hidden units have similar parameter counts with Mamba with 16 (32) hidden units.
GRU1 LSTM1 Mamba1
Layer Num 1 1 1
Hidden Size 32 32 16
Params 6529 8641 9679
$T=1$ Memory 2397MiB 2507MiB 2157MiB
Time 59min 58min 95min
Rel. $L_{2}$ Err 3.909E-2$\pm$9.204E-3 3.720E-2$\pm$1.684E-2 3.991E-2$\pm$8.022E-3
$T=2$ Memory 3229MiB 3447MiB 2919MiB
Time 117min 94min 193min
Rel. $L_{2}$ Err NaN NaN 1.338E-1$\pm$5.959E-3
$T=3$ Memory 4063MiB 4389MiB 3687MiB
Time 181min 134min 296min
Rel. $L_{2}$ Err NaN NaN 2.470E-1$\pm$1.117E-2
GRU2 LSTM2 Mamba2
Layer Num 1 1 1
Hidden Size 64 64 32
Params 25345 33665 22593
$T=1$ Memory 3039MiB 3257MiB 2265MiB
75min 61min 116min
NaN 6.158E-1$\pm$7.201E-2 3.417E-2$\pm$4.156E-3
$T=2$ Memory 4511MiB 4949MiB 3143MiB
146min 114min 230min
NaN NaN 1.230E-1$\pm$3.954E-3
$T=3$ Memory 5983MiB 6637MiB 4033MiB
215min 169min 346min
NaN NaN 2.157E-1$\pm$1.643E-2
@@
Results are shown in Table 9. In the chaotic Lorenz system with a terminal time equal to $T=2$ or $T=3$, we find that LSTM and GRU models of various sizes diverge to not-a-number (NaN) values during training due to their model instability. Chaotic systems are known for their instability and sensitivity to minor changes in their initial values. A minor difference in initial values will lead to a considerable difference in ODE trajectories, especially during long-time integration. This, in turn, implies that the model has to be able to capture the subtle differences in initial values and react differently, which is a challenging task. Hence, we observe that LSTM and GRU converge well in the $T=1$ case where the time is short, and the chaotic dynamics have not developed yet. However, as we reach longer times, LSTM and GRU diverge due to the chaotic dynamics, even though we train them using a decaying learning rate for stabilization. This reflects RNNs’ inherent instability in lengthy sequence modeling (length = 1K, 2K, and 3K here) due to unstable gradient explosion even for its modern variants, namely LSTM and GRU [2, 39]. In contrast, Mamba always converges in all systems, demonstrating its superior stability and robustness regardless of the long-time integration.
@@ 4.7 Quantifying the Extrapolation Error
We have extensively tested Mamba’s extraordinary out-of-distribution (OOD) extrapolation capability in (1) OOD operator learning following Cao et al. [4] in Section 4.3, (2) extrapolation to long-time integration problems while only training on shorter time 4.5, (3) extrapolation to treatment schedule parameters in the PK-PD model in Section 4.8.4. In this section, we follow Zhu et al. [56] to quantify further various models’ extrapolation errors under different train and test input function distributions.
Regarding the ODE operators, we test the antiderivative and the gravity pendulum operators in Section 4.1. Following Zhu et al. [56], the input functions $u(t)$ are randomly drawn from a mean zero Gaussian random field (GRF) with a length scale parameter $l$, i.e., the input functions follow the distribution $\mathcal{GP}\left(0,k_{l}(x,y)\right)$ where $k_{l}(x,y)=\exp\left(-\frac{(x-y)^{2}}{2l^{2}}\right)$ is the RBF kernel with length scale $l$ and $\mathcal{GP}$ denotes a Gaussian process. A smaller correlation length scale $l$ leads to less smooth functions, while a larger correlation length scale $l$ leads to smoother functions.
@@
We generate 10K training and validation data under $l_{train}$ and 10K testing data under each $l_{test}$. The validation is conducted under an IID set following the same $l_{train}$. In real-world applications, we cannot access the testing distribution in advance and must use the same source for validation. We keep all settings, e.g., implementation details, training, and validation procedures, the same as those in Section 4.1 except the train and test input function distributions. We finally obtain the best hyperparameters and model structures the same as in Section 4.1.
@@
The detailed quantitative results are shown in Table 13 (gravity pendulum Ex-), Table 14 (gravity pendulum Ex+), Table 15 (anti-derivative Ex-), Table 16 (anti-derivative Ex+) in B. Their visualization is shown in Figure 11 (gravity pendulum Ex-), Figure 12 (gravity pendulum Ex+), Figure 13 (anti-derivative Ex-), Figure 14 (anti-derivative Ex+). Each figure shows the visualized results for one of the four settings, gravity pendulum or anti-derivative operator in the Ex- or the Ex+ setting. Furthermore, in each figure, we show Mamba’s comparison with RNNs, including GRU and LSTM, in subfigure (a), with neural operators (DeepONet, FNO, and LNO) in (b), and with transformers (Transformer, Oformers, and GNOT) in (c).
Regarding the pendulum problem, Mamba is the best in both Ex- problems (Figure 11 and Table 13) and Ex+ (Figure 12 and Table 14) in general. More concretely, in the Ex- setting (Figure 11 and Table 13), Mamba’s error remains stable under the test distribution shift. Although certain transformers (Oformer-G and Oformer-F) also exhibit stability against test distribution shift, their error is larger than Mamba. Other models, including RNNs and neural operators, exhibit increases in errors when the test distribution shift is more significant, i.e., their error grows larger as $l_{test}=0.1$ gradually moves to $l_{test}=1$. Meanwhile, regarding the Ex+ setting of gravity pendulum (Figure 12 and Table 14), all models’ errors grow with test distribution shift as $l_{test}=1$ gradually moves to $l_{test}=0.1$ where the training set length scale is $l_{train}=1$. However, Mamba exhibits the slowest error growth and consistently performs the best.
Regarding the anti-derivative operator, Mamba is the best in the Ex- setting (Figure 13 and Table 15), and its Ex- error remains stable. We also observe that Oformers [28] can exhibit stability against test distribution shift but perform worse than Mamba in general. Moreover, other models, including RNNs and neural operators, exhibit increases in errors when the test distribution shift is more significant, i.e., their error grows larger as $l_{test}=0.1$ gradually moves to $l_{test}=1$. This is similar to the Ex- case in the gravity pendulum case. Regarding the anti-derivative operator’s Ex+ setting (Figure 14 and Table 16), Mamba is good in the majority of test cases. Mamba is only beaten in the case with $l_{train}=1$ and $l_{test}=0.1$, although not by much, and it is close to the best model. Quantitatively, Mamba’s relative $L_{2}$ error for the anti-derivative operator’s Ex+ setting with $l_{train}=1$ and $l_{test}=0.1$ is 3.091E-2, while the best model LNO’s is 1.721E-2.
@@ 4.8.1 Model and Data Generation
Simeoni et al. [48] introduced a streamlined PK-PD model to simplify the complex dynamics of tumor growth, making it more applicable for the preclinical development of oncology drugs. This model, formulated with a system of ODEs, connects the drug dosing regimen to the tumor growth dynamics observed in animal studies. The model differentiates between untreated and treated animals; in untreated animals, tumor growth follows an initial exponential phase and then a linear phase, whereas, in treated animals, the growth rate decreases in relation to the drug concentration and the number of proliferating tumor cells. The differential equation for the unperturbed growth of the tumor mass $w(t)$ is given by:
@@
In this work, we aim to learn the mapping from $I_{1}(t)$ in the PK model, representing the drug dose and administration schedule in a multi-dose regimen, to the patient-specific function $w(t)$, which reflects tumor weight based on individual patient PD model parameters $k_{1}$ and $k_{2}$. These parameters indicate disparate responses to treatment across different species.
@@ 4.8.2 Learning the PK-PD Model’s Operator
We generated a dataset containing 5000 (5K) species for both the PK and PD models based on the procedure in Section 4.8.1. The dataset comprises the dose amount and dose schedule as inputs (represented as $I_{1}(t)$ in the PK model), the species-specific PD model parameters $k_{1}$ and $k_{2}$, as well as $w(t)$, which is the model output. Upon feeding the model with the values of $k_{1}$, $k_{2}$, and $I_{1}(t)$, the resulting output is the tumor size in grams. It is observed that different dynamics and tumor sizes are exhibited by different species or treatment schedules. We divide the dataset into 4K training and 1K testing. During validation and model hyperparameter selection, we leave 20% of the training data as validation and train on the rest 80%. After the model hyperparameter selection, we returned the validation data to training and optimized the model with a total of 4K training data. We append the temporal grid to the input data tensor for augmentation.
@@
Here are the implementation details. The training loss is MSE, and the test metric is the relative $L_{2}$ error. We train all models for 10,000 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [27] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training. We keep the number of parameters of different models similar for a fair comparison. We run the code for five independent random seeds and report the average.
@@
The model structure hyperparameter settings are summarized below. Similar to Section 4.1, we set a parameter count budget of approximately 10,000. For all time series models, we search the model depth in $1\sim 3$ and choose the corresponding hidden sizes to satisfy the budget. We found that shallow one-layer and wide models are better than deeper ones in speed and accuracy. LSTM and GRU are one-layer models with 32 hidden dimensions. Regarding Mamba, we opt for a block/layer with skip connection and layer normalization, with 16 model dimensions and 16 intermediate dimensions. We opt for a one-layer Transformer with 40 hidden and feedforward dimensions, and four heads. For all three Oformers with vanilla, Galerkin, or Fourier attention, we opt for one-layer Oformers with 24 hidden and feedforward dimensions and four heads. GNOT has one layer, 16 hidden feedforward dimensions, four heads, and two experts. We choose the same one-layer comparable structures for all time series models for a fair comparison. Regarding NOs, we keep DeepONet’s trunk and branch nets as MLP with the same structure, whose depth is $2\sim 5$. We find the deepest model is the best, i.e., DeepONet’s trunk and branch nets are all five-layer networks with 24 hidden dimensions. We keep the same structure as in Section 4.1, but since the input sequence is longer, DeepONet here has a larger parameter count. We searched for the best FNO structure among $1\sim 4$ layers and $8$ modes and obtained the best four-layer model with 16 hidden dimensions and eight modes. We searched for the best LNO models with $8/16$ hidden dimensions and $4/8$ modes, and finally, we chose a one-layer LNO with eight hidden dimensions and four modes and another one-layer LNO with 16 hidden dimensions and eight modes. We also tried larger LNOs, but this small LNO model performs the best on the validation set. We report both LNO performances.
Figure 15: Visualization of various models’ loss trajectories for the PK/PD model with IID train-test data corresponding to Section 4.8.2 in this paper. The detailed quantitative results for this setting are presented in Table 10. We visualize some inputs and Mamba’s predictions in Figure 16. We visualize Mamba’s relative $L_{2}$ error concerning time in Figure 17. Subfigure (a) is Mamba versus RNNs; subfigure (b) is Mamba versus NOs; and subfigure (c) is Mamba versus Transformers.
Model Params Memory Time Relative $L_{2}$ Error
LSTM 8641 1985MiB 16min 2.117E-3$\pm$1.512E-4
GRU 6529 1925MiB 16min 2.125E-3$\pm$1.201E-4
DeepONet 38568 1425MiB 15min 4.732E-1$\pm$1.169E-1
FNO 9377 2021MiB 30min 2.917E-3$\pm$4.158E-5
LNO (Large) 6769 1991MiB 24min 8.269E-1$\pm$5.178E-2
LNO (Small) 1913 1935MiB 22min 5.295E-3$\pm$1.063E-3
Transformer 10321 2615MiB 56min 2.847E-2$\pm$5.340E-3
Oformer-V 11281 2469MiB 49min 7.708E-3$\pm$3.901E-4
Oformer-G 11233 1591MiB 46min 9.101E-3$\pm$7.357E-4
Oformer-F 11233 1591MiB 46min 8.129E-3$\pm$6.255E-4
GNOT 11235 1649MiB 91min 5.437E-3$\pm$9.589E-4
Mamba 9697 1707MiB 19min 1.685E-3$\pm$2.929E-4
@@
Table 10: Results for the PK-PD model corresponding to Section 4.8.2. The best model is bold, the second best is red, and the third best is blue. We visualize the loss trajectory in Figure 15. We visualize some inputs and Mamba’s predictions in Figure 16. We visualize Mamba’s relative $L_{2}$ error concerning time in Figure 17.
@@
Figure 16: Visualization of Mamba’s prediction in PK-PD model corresponding to Section 4.8.2. We plot the input values and the outputs for test data. (a) and (b): input/output (io) for test data No. 0; (c) and (d): io for test data No. 100; (e) and (f): io for test data No. 200; (g) and (h): io for test data No. 300. Note that the heights and positions of the input spikes vary for different data sets, and the PD parameters differ based on each individual’s response to treatment. Moreover, we visualize loss trajectory in Figure 15 and Mamba’s relative $L_{2}$ error with respect to time in Figure 17. The full quantitative results are shown in Table 10.
@@
Results are shown in Table 10. Mamba continues to be the best-performing model among multiple strong baselines. We visualize the inputs and outputs of this PK-PD model in Figure 16 and plot Mamba’s relative $L_{2}$ error with respect to time in Figure 17. The results show that Mamba can capture the PK-PD model’s operator containing discontinuous inputs. The Mamba model’s relative $L_{2}$ error of 1.685E-3 is equivalent to an error of 0.1685% of the tumor size. Thus, for a tumor volume of 10 grams, the resulting error is 16.85 milligrams. Despite a limited 10,000 parameter count budget for Mamba, it can already achieve a satisfactorily low error without the need to further expand the model size for better results. Lastly, we visualize different models’ training loss trajectories in Figure 15. Mamba maintains its advantage in faster convergence speed and better optimization results throughout the optimization.